# Global Non-Axisymmetric Hall Instabilities in a Rotating Plasma

### v1

@@ 2.1 Local dispersion relation
Equation (14) implies that decreasing rotation profiles, $\Gamma<0$, with sufficiently strong shear are locally unstable. The first term on the right-hand side is simply the local frequency of a whistler wave. The other terms arise from the variable transformation from $\psi$ to $u$, and have the potential to be either stabilizing or destabilizing. The final two terms represent the effect of magnetic shear from the radially dependent azimuthal field. Note that for weak magnetic fields, $|\mathrm{H}|\rightarrow 0$, the instability criterion can be satisfied by modest shear. Some aspects of this local analysis are qualitatively similar to the behavior of the local $m=0$ Hall-MRI dispersion relations (M. Wardle, 1999; S. A. Balbus & C. Terquem, 2001). For comparison, the $m=0$ limit of (14) is
@@
Here we have used the simplification $\mathrm{H}\propto kd_{i}B_{z}$ when $m=0$. The third term on the right-hand side is essentially the rotational transform of the magnetic field, $\iota$. The expression in equation (15) is similar but not identical to the necessary criterion derived by Balbus and Terqueum S. A. Balbus & C. Terquem (2001) in the axisymmetric Hall-MHD case without the stabilizing contribution from magnetic field-line bending proportional to $(\boldsymbol{k}\cdot\boldsymbol{V}_{A})^{2}$. It is also similar to the criterion given by Kunz M. W. Kunz (2008) for the HSI in a non-rotating sheared flow using the local shearing box approximation. The disappearance of the field-line bending from equation (15) compared to equation (85) of S. A. Balbus & C. Terquem (2001) or equation (42) of M. W. Kunz (2008) is a direct result of the EMHD approximation. The effect of magnetic shear is either stabilizing or destabilizing depending on the relative signs of $m$ and $\mathrm{H}_{\phi}/\mathrm{H}$. If the field is purely azimuthal, $\mathrm{H}_{\phi}/\mathrm{H}\propto 1/m$, and the contribution is stabilizing. In the general case of $B_{\phi}\neq 0,B_{z}\neq 0$, there could exist a location where $m/s^{2}|\boldsymbol{k}|^{2}+\mathrm{H}_{\phi}/s\mathrm{H}=0$, minimizing the stabilizing contribution of that term.
@@
Let us evaluate some limiting cases of equations (14) and (16). As $k\rightarrow\infty$, we have $\mathrm{H}\sim kd_{i}V_{A,z}/r_{1}\Omega_{0}$, $\gamma\sim\sqrt{-k\Gamma\mathrm{H}-(k^{2}+(B_{\phi}^{2}/B_{z}^{2})/s^{2})\mathrm{H}^{2}}$. Thus, in general, large $k$ modes are only unstable for sufficiently small $|\mathrm{H}|$. Interestingly, nonzero $B_{\phi}$ will stabilize the high $k$ modes even as $B_{z}\rightarrow 0$ if $kd_{i}V_{A,\phi}/r_{1}\Omega_{0}>s\Gamma/4$. This stabilization of the high $k$ modes is independent of any dissipative mechanism (e.g. resistivity or viscosity). Also, if $|\mathrm{H}|$ is fixed, there are no growing modes for $k^{2}>r_{1}\Omega_{0}\Gamma/d_{i}V_{A,z}$. When $m\rightarrow\infty$, equation (14) gives $k\Gamma/\mathrm{H}>m^{2}/s^{2}$, where $\mathrm{H}\rightarrow(m/s)d_{i}V_{A,\phi}/r_{1}\Omega_{0}$. The high-$m$ modes will certainly be stable for $|m|\gg|sk\Gamma r_{1}\Omega_{0}/d_{i}V_{A,\phi}|$. For a given field strength, short wavelength whistler waves are stable.
@@
The local criterion for instability in equation (14) and subsequent expression for the growth rate (equation (16)) provide physical intuition, but extracting quantitative results from these expressions presents some difficulty. There is an ambiguity both in the choice of the radial wavenumber, $\delta$, as well as the radial location at which the various $s$-dependent terms should be evaluated. Moreover, we observe in numerical calculations that unstable eigenmodes can extend throughout the domain – invalidating assumptions of locality. It is also important to note that a failure to meet the local instability criterion in equation (14) does not necessarily preclude the existence of global unstable eigenmodes. In particular, when $k=0,m\neq 0$ in the azimuthal magnetic field case, there are unstable global modes associated with whistler waves being amplified by the flow. This specific example is discussed further at the end of section 2.2.
2.2 Global Eigenvalue Problem
@@
Equation (18) also defines a variational problem for $\omega$. Namely, $\omega$ is considered as a functional of $u$ given fixed $m,k,\Omega,\mathrm{H}$ and $\mathrm{H}_{\phi}$, then the vanishing of the first variation of $\omega$ is equivalent to $u$ being a solution of equation (12). The discriminant, $\Delta\equiv\mathcal{B}^{2}-4\mathcal{A}\mathcal{C}$ can also be viewed as a functional of $u$. The minimizer, $v$, of $\Delta$ subject to the constraint that $v$ solves the original ODE, if it exists, should correspond to the most unstable eigenfunction of the problem when $\Delta<0$. If the minimum value of $\Delta$ is positive, then the system is stable.
@@
Figure 2: Numerically computed growth rates and frequencies of the fastest growing radial mode in a disc with vertical magnetic field in the electron MHD limit with $m=1,k=\pi/4$ as a function of $\mathrm{H}\propto kB_{z}$ at several aspect ratios.
2.3 Numerical Results for EMHD
We present numerical results from the electron MHD system using a spectral element method to reduce the radial differential eigenvalue problem to a generalized linear eigenvalue problem that is solved using an appropriate LAPACK routine. In terms of the normalized radial coordinate, the Keplerian rotation profile is $\Omega\propto s^{-3/2}$, and the normalized rotation shear in equation (12) is $\Gamma(s)\propto-s^{3/2}$.
@@ 2.3 Numerical Results for EMHD
Here, we consider the most global non-axisymmetric mode, i.e. $m=1,k=2\pi r_{1}/\ell$. This value of $k$ corresponds to the minimum non-zero normalized vertical wavenumber that satisfies periodic boundary conditions at the ends of the cylinder. In fig. 2, we plot the growth rate of the fastest growing numerically computed mode as a function of $\mathrm{H}$ at several values of the aspect ratio, $A$, defined as $A\equiv\ell/(r_{2}-r_{1})$. The variation in aspect ratio is achieved by fixing $\ell/r_{1}=8\implies k=\pi/4$ and adjusting the distance between the inner and outer boundary, $r_{2}/r_{1}$. As the aspect ratio (i.e. disk thickness) increases, the outer boundary is moved closer to the inner boundary, and the range of $\mathrm{H}$ that supports unstable modes decreases. In the thin disk limit limit $A\ll 1$, the outer boundary is far enough away that it does not affect the fastest growing mode. Fig. 3 compares plots of the numerically computed radial eigenfunction $\psi$ of equation (11) a low aspect ratio ($A=0.5$) and a high aspect ratio ($A=8$) case. In the high aspect ratio case, the unstable mode is localized near the inner boundary, and is substantially damped to the right of the co-rotation point. In the low aspect ratio case, the co-rotation point is closer to the outer boundary, and the mode has nontrivial amplitude throughout the domain.
@@
Figure 3: Plots of the eigenfunction $\psi$ associated with the fastest growing modes from the vertical field case at $A=0.5$, $\mathrm{H}=0.08$ (left), and $A=8$, $\mathrm{H}=0.08$ (right). The dashed red lines denote the point $\bar{\omega}=0$ for each mode. Note the difference in radial scales.
@@
Turning to the disk configuration with purely azimuthal magnetic field, we note two unique feature that are absent in the uniform axial field case. Since $\boldsymbol{B}\parallel\hat{\boldsymbol{\phi}}$, the propagation direction of whistler waves is also along the direction of the bulk flow. Also, the magnetic field strength varies with radial location in accordance with the current-free condition $(rB_{\phi})^{\prime}=0$. In this case, there is also a transition from the low-field, local modes to higher field global modes that depends on the aspect ratio. Fig. 4 plots the growth rates and frequencies of the $m=1,k=\pi/4$ mode
@@
Figure 4: Numerically computed growth rates and frequencies of the fastest growing radial mode in a disc with azimuthal magnetic field in the electron MHD limit with $m=1,k=\pi/4$ as a function of $\mathrm{H}\propto B_{\phi}$ evaluated at $s=1$ at several aspect ratios.
@@
Since the location of the outer boundary does not appear explicitly in in eqs. (14) or (16), one would not expect to be able to explain this behavior based on local theory alone. Qualitatively, we can deduce some effect from the local expression by introducing a heuristic expression for the radial wavenumber $\delta\propto 1/(r_{2}/r_{1}-1)$. With this expression for $\delta$, decreasing aspect ratio yields a smaller $\delta^{2}$. For a fixed azimuthal mode number, $m$, the minimum value of $m^{2}/r^{2}$ attained in the domain also decreases with decreasing aspect ratio. These two effects result in a correspondingly smaller value for $|\boldsymbol{k}|$ appearing on the right-hand side of equation (16), suggesting larger local growth rates.
2.4 Over-reflection of $k=0$ modes in azimuthal field
@@
When $B_{\phi}\neq 0$, the EMHD model has unstable solutions that correspond to unstable modes with $k=0$ and $m\neq 0$. Instability in this case is an interesting finding for two reasons. First, since the term that contains the flow shear, $k\Gamma/\mathrm{H}$, vanishes when $k=0$, one may be under the impression that the instability drive is absent. It is, however, not explicitly required that $k\Gamma/\mathrm{H}\neq 0$ for instability to occur. Second, when $k=0$ the local criterion in equation (14) predicts stability for $m^{2}+s^{2}\delta^{2}>1/4$. Since the minimum nonzero value of $m^{2}$ permitted by the cylindrical symmetry of the problem is $m^{2}=1$, any choice of $\delta\in\mathbb{R}$ is stable. The existence of unstable modes with $m=1,k=0$ highlights the potential danger of drawing stability conclusions based solely on local WKB analysis.
@@
Using the global criteria based of equation (17), we find that $\gamma\neq 0$ is possible if $\int_{1}^{r_{2}/r_{1}}(\omega_{r}-m\Omega(s))s^{4}|u|^{2}ds=0$. The important condition in this case is whether the rotation profile varies sufficiently so that $\omega_{r}-m\Omega(r)$ changes signs in the domain. This requirement for instability suggests that the unstable modes are whistler waves being destabilized by a ‘corotation amplifier’ mechanism (R. Narayan et al., 1987; D. Tsang & D. Lai, 2008).
@@
Figure 5: Left: Plot of the square of the effective refractive index for whistler waves, $Q(\omega,s)$ for the fastest growing $m=1,k=0$ mode with $d_{i}V_{A,\phi}=0.58$. The shaded region is where $Q(\omega,s)<0$, and the zeros of $Q$ are where $\bar{\omega}^{2}=|\boldsymbol{k}|^{2}d_{i}^{2}\omega_{A}^{2}$ The real and imaginary parts of $\psi$ are also plotted. Right: plot of the growth rate for $m=1,k=0$ modes in the EMHD model as a function of $\mathrm{H}_{\phi}$ for aspect ration $A=0.5$ along with analytic estimate based on asymptotics equation (A17).

### v2

@@ III.1 Local dispersion relation
Equation (21) implies that decreasing rotation profiles, $\Gamma<0$, with sufficiently strong shear are locally unstable. The first term on the right-hand side is simply the local frequency of a whistler wave. The other terms arise from the variable transformation from $\psi$ to $u$, and have the potential to be either stabilizing or destabilizing. The final two terms represent the effect of magnetic shear from the radially dependent azimuthal field. Note that for weak magnetic fields, $|\mathrm{H}|\rightarrow 0$, the instability criterion can be satisfied by modest shear. Some aspects of this local analysis are qualitatively similar to the behavior of the local $m=0$ Hall-MRI dispersion relations (Wardle, 1999; Balbus and Terquem, 2001). For comparison, the $m=0$ limit of (21) is
@@
Here we have used the simplification $\mathrm{H}\propto kd_{i}B_{z}$ when $m=0$. The fourth term on the right-hand side is essentially the rotational transform, or field-line twist, of the magnetic field, $\iota$. The expression in equation (22) is similar but not identical to the necessary criterion derived by Balbus and Terquem (2001) in the axisymmetric Hall-MHD case without the stabilizing contribution from magnetic field-line bending proportional to $(\boldsymbol{k}\cdot\boldsymbol{V}_{A})^{2}$. It is also similar to the criterion given by Kunz Kunz (2008) for the HSI in a non-rotating sheared flow using the local shearing box approximation. The disappearance of the field-line bending term from equation (22) compared to equation (85) of Balbus and Terquem (2001) or equation (42) of Kunz (2008) is a direct result of the EMHD approximation. The effect of magnetic shear is either stabilizing or destabilizing depending on the relative signs of $m$ and $\mathrm{H}_{\phi}/\mathrm{H}$. If the field is purely azimuthal, $\mathrm{H}_{\phi}/\mathrm{H}\propto 1/m$, and the contribution is stabilizing. In the general case of $B_{\phi}\neq 0,B_{z}\neq 0$, there could exist a location where $m/s^{2}|\boldsymbol{k}|^{2}+\mathrm{H}_{\phi}/s\mathrm{H}=0$, minimizing the stabilizing contribution of that term.
@@
Let us evaluate some limiting cases of equations (21) and (23). As $k\rightarrow\infty$, we have $\mathrm{H}\sim kd_{i}V_{A,z}/r_{1}\Omega_{0}$, $\gamma\sim\sqrt{-k\Gamma\mathrm{H}-(k^{2}+(B_{\phi}^{2}/B_{z}^{2})/s^{2})\mathrm{H}^{2}}$. Thus, in general, large $k$ modes are only unstable for sufficiently small $|\mathrm{H}|$. Interestingly, nonzero $B_{\phi}$ will stabilize the high $k$ modes even as $B_{z}\rightarrow 0$ if $kd_{i}V_{A,\phi}/r_{1}\Omega_{0}>s\Gamma/4$. This stabilization of the high $k$ modes is independent of any dissipative mechanism (e.g. resistivity or viscosity). Also, if $|\mathrm{H}|$ is fixed, there are no growing modes for $k^{2}>r_{1}\Omega_{0}\Gamma/d_{i}V_{A,z}$. When $m\rightarrow\infty$, equation (21) gives $k\Gamma/\mathrm{H}>m^{2}/s^{2}$, where $\mathrm{H}\rightarrow(m/s)d_{i}V_{A,\phi}/r_{1}^{2}\Omega_{0}$. The high-$m$ modes will certainly be stable for $|m|\gg|sk\Gamma r_{1}\Omega_{0}/d_{i}V_{A,\phi}|$. For a given field strength, short wavelength whistler waves are stable.
@@
The local criterion for instability in equation (21) and subsequent expression for the growth rate (equation (23)) provide physical intuition, but extracting quantitative results from these expressions presents some difficulty. There is an ambiguity both in the choice of the radial wavenumber, $\delta$, as well as the radial location at which the various $s$-dependent terms should be evaluated. Moreover, we observe in numerical calculations that unstable eigenmodes can extend throughout the domain – invalidating assumptions of locality. It is also important to note that a failure to meet the local instability criterion in equation (21) does not necessarily preclude the existence of global unstable eigenmodes. In particular, when $k=0,m\neq 0$ in the azimuthal magnetic field case, there are unstable global modes associated with whistler waves being amplified by the flow. This specific example is discussed further at the end of Section III.4.
III.2 Global Eigenvalue Problem
@@
Figure 13: Numerically computed growth rates and frequencies of the fastest growing radial mode in a disk with axial magnetic field in the electron MHD limit with $m=1,k=\pi/4$ as a function of $\mathrm{H}\propto kB_{z}$ at several aspect ratios.
III.3 Numerical Results for EMHD
We present numerical results from the electron MHD system using another version of the CYL_SPEC code described in Section II.1
@@ III.3 Numerical Results for EMHD
Here, we consider the most global non-axisymmetric mode, i.e. $m=1,k=2\pi r_{1}/\ell$. This value of $k$ corresponds to the minimum non-zero normalized axial wavenumber that satisfies periodic boundary conditions at the ends of the cylinder. In figure 13, we plot the growth rate of the fastest growing numerically computed mode as a function of $\mathrm{H}$ at several values of the aspect ratio, $A$, defined as $A\equiv\ell/(r_{2}-r_{1})$. The variation in aspect ratio is achieved by fixing $\ell/r_{1}=8\implies k=\pi/4$ and adjusting the distance between the inner and outer boundary, $r_{2}/r_{1}$. As the aspect ratio (i.e. disk thickness) increases, the outer boundary is moved closer to the inner boundary, and the range of $\mathrm{H}$ that supports unstable modes decreases. In the thin disk limit limit $A\ll 1$, the outer boundary is far enough away that it does not affect the fastest growing mode. Figure 14 compares plots of the numerically computed radial eigenfunction $\psi$ of equation (18) a low aspect ratio ($A=0.5$) and a high aspect ratio ($A=8$) case. In the high aspect ratio case, the unstable mode is localized near the inner boundary, and is substantially damped to the right of the co-rotation point. In the low aspect ratio case, the co-rotation point is closer to the outer boundary, and the mode has nontrivial amplitude throughout the domain.
@@
Figure 14: Plots of the eigenfunction $\psi$ associated with the fastest growing modes from the axial field case at $A=0.5$, $\mathrm{H}=0.08$ (left), and $A=8$, $\mathrm{H}=0.08$ (right). The dashed red lines denote the point $\bar{\omega}=0$ for each mode. Note the difference in radial scales.
@@
Turning to the disk configuration with purely azimuthal magnetic field, we note two unique feature that are absent in the uniform axial field case. Since $\boldsymbol{B}\parallel\hat{\boldsymbol{\phi}}$, the propagation direction of whistler waves is also along the direction of the bulk flow. Also, the magnetic field strength varies with radial location in accordance with the current-free condition $(rB_{\phi})^{\prime}=0$. In this case, there is also a transition from the low-field, local modes to higher field global modes that depends on the aspect ratio. Figure 15 plots the growth rates and frequencies of the $m=1,k=\pi/4$ mode
@@
Figure 15: Numerically computed growth rates and frequencies of the fastest growing radial mode in a disk with azimuthal magnetic field in the electron MHD limit with $m=1,k=\pi/4$ as a function of $\mathrm{H}\propto B_{\phi}$ evaluated at $s=1$ at several aspect ratios.
@@
Since the location of the outer boundary does not appear explicitly in in eqs. (21) or (23), one would not expect to be able to explain this behavior based on local theory alone. Qualitatively, one can infer some effect by introducing a heuristic expression for the radial wavenumber $\delta\propto 1/(r_{2}/r_{1}-1)$ into the local dispersion relation, equation (23). With this expression for $\delta$, decreasing aspect ratio yields a smaller $\delta^{2}$. For a fixed azimuthal mode number, $m$, the minimum value of $m^{2}/r^{2}$ attained in the domain also decreases with decreasing aspect ratio. These two effects result in a correspondingly smaller value for $|\boldsymbol{k}|$ appearing on the right-hand side of equation (23), suggesting larger local growth rates.
III.4 Over-reflection of $k=0$ modes in azimuthal field
@@
When $B_{\phi}\neq 0$, the EMHD model has unstable solutions that correspond to unstable modes with $k=0$ and $m\neq 0$. Instability in this case is an interesting finding for two reasons. First, since the term that contains the flow shear, $k\Gamma/\mathrm{H}$, vanishes when $k=0$, one may be under the impression that the instability drive is absent. It is, however, not explicitly required that $k\Gamma/\mathrm{H}\neq 0$ for instability to occur. Second, when $k=0$ the local criterion in equation (21) predicts stability for $m^{2}+s^{2}\delta^{2}>1/4$. Since eigenmodes in the cylindrical domain are $2\pi$-periodic in $\phi$, the minimum nonzero value of $m^{2}$ permitted is $m^{2}=1$. Therefore, any choice of $\delta\in\mathbb{R}$ is stable. The existence of unstable modes with $m=1,k=0$ highlights the potential danger of drawing stability conclusions based solely on local WKB analysis.
@@
Using the global criteria based of equation (24), we find that $\gamma\neq 0$ is possible if $\int_{1}^{r_{2}/r_{1}}(\omega_{r}-m\Omega(s))s^{4}|u|^{2}ds=0$. The important condition in this case is whether the rotation profile varies sufficiently so that $\omega_{r}-m\Omega(r)$ changes signs in the domain. This requirement for instability suggests that the unstable modes are whistler waves being destabilized by a ‘corotation amplifier’ mechanism (Narayan et al., 1987; Tsang and Lai, 2008).
@@
Figure 16: Left: Plot of the square of the effective refractive index for whistler waves, $Q(\omega,s)$ for the fastest growing $m=1,k=0$ mode with $d_{i}V_{A,\phi}=0.58$. The shaded region is where $Q(\omega,s)<0$, and the zeros of $Q$ are where $\bar{\omega}^{2}=|\boldsymbol{k}|^{2}d_{i}^{2}\omega_{A}^{2}$ The real and imaginary parts of $\psi$ are also plotted. Right: plot of the growth rate for $m=1,k=0$ modes in the EMHD model as a function of $\mathrm{H}_{\phi}$ for aspect ration $A=0.5$ along with analytic estimate based on equation (A17).